On a given day, each computer in the lab has at most one crash. There is a 5% chance
that a computer has a crash during the day, independent of the performance of any
other computers in the lab. Find the probability that on any given day, there are;
(a) exactly 3 crashes
(b) at most 3 crashes
(c) at least 3 crashes
(d) more than 3 and less than 6
1
Expert's answer
2020-04-28T15:16:12-0400

Assume n - number of computers,k - number of crashed computers according to a bernoulli formula:

a)Pn(k=3)=C(n,3)*0.053*0.95(n-3)

b)Pn(k"\leq"3)=Pn(0)+Pn(1)+Pn(2)+Pn(3)=C(n,0)*0.050*0.95n+C(n,1)*0.051*0.95n-1+C(n,2)*0.052*0.95n-2+C(n,3)*0.053*0.95n-3

c)Pn(k"\geq" 3)=Pn(3)+....+Pn(n)=C(n,3)*0.053*0.95n-3+...+C(n,n)*0.05n

d)Pn(3<k<6)=Pn(4)+Pn(5)=C(n,4)*0.054*0.95n-4+C(n,5)*0.055*0.95n-5